OpenStudy (anonymous):

solve for t 1.75 t - 4.5 = 7.75

OpenStudy (anonymous):

1.75t = 12.25 t = 7

OpenStudy (anonymous):

i got 1.8 -.-

OpenStudy (anonymous):

rewrite your 1.75t as 1 and 3/4 T = 12.25

OpenStudy (anonymous):

so 7/4T = 12.25

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

well 1.75T is like saying 1 and 3 quarters of T, correct?

OpenStudy (anonymous):

yes, but dont do fractions please, confuses me and we are supposed to write it out.

OpenStudy (anonymous):

Well first thing's first, get T alone.

OpenStudy (anonymous):

1.75T technically.

OpenStudy (anonymous):

okay. when i did it 1.75t - 4.5 = 7.75 1.75t - 4.5 - 4.5 = 7.75 - 4.5 1.75t =3.25 t=1.8

OpenStudy (anonymous):

you don't want to subtract -4.5 Think of it this way. you're doing 1.75t - 4.5 - 4.5 = 7.75 - 4.5 which gives you 1.75t - 9 = 3.25 see where you're going wrong?

OpenStudy (anonymous):

did you mean to put 1.75 - 4.5 + 4.5? how would you get 9 ?

OpenStudy (anonymous):

You subtracted -4.5 from -4.5 -4.5 minus 4.5 = -9

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

-4.5 does not cancel out - 4.5. positive 4.5 cancels out - 4.5.

OpenStudy (anonymous):

i see. so how should i go about this?

OpenStudy (anonymous):

so what you want to do is 1.75t -4.5 + 4.5 = 7.75 + 4.5 so that the -4.5 and 4.5 add to be equal to 0 and you are left with 1.75t = 12.25

OpenStudy (anonymous):

i just had my operations wrong, i get it.

OpenStudy (anonymous):

mmhmm, you are rather smart at this stuff; you just get mixed up on your signs a lot is all :)

OpenStudy (anonymous):

i think too much, thanks /:

OpenStudy (anonymous):

so my answer is 1.75t + 12.25 or 7?

OpenStudy (anonymous):

well you're not at the answer yet. after you remove the 4.5 you are left with 1.75t = 12.25 how would you get t to be not multiplied by 1.75?

OpenStudy (anonymous):

divide it by 1.75 and divide 12.25 by 1.75 to get 7?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

┌─┐ ┴─┴ ಠ_ರೃ until my next problem.